OpenStudy (anonymous):

If the arc length shown in blue is 15.5 inches, then theta to the nearest hundredth of a radian is _____.

Parth (parthkohli):

\[r\times \theta = \rm arc~length \]

OpenStudy (anonymous):

So it would be, \[12*\theta=15.5\] And we have to find \[\theta\] so, \[15.5*12=\theta\] Am I doing it right?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (***[isuru]***):

yep... ur steps r correct.. :)

OpenStudy (anonymous):

okay, thank you!

OpenStudy (***[isuru]***):

u r welcome!